OpenStudy (anonymous):

help : multiply out (9-6i)^2

OpenStudy (anonymous):

(9 - 6i)(9 - 6i) you foil 81 - 54i - 54i - 36 (-6i)(-6i) is -36 because - times a - is a positive but i^2 = -1 add like terms

OpenStudy (anonymous):

81 - 36 = 45 -54i - 54i = -108i so you get 45 - 108i

OpenStudy (anonymous):

y (-6i)(-6i)=-36 , it's not 36i^2?

OpenStudy (anonymous):

i^2 = -1 by definition so 36(-1) = -36

OpenStudy (anonymous):

thank u.